DogGirl27:

A rectangle has an area of 40 square units. The length is 6 units greater than the width. What are the dimensions of the rectangle? 8 by 5 10 by 4 11 by 9 13 by 7

Alex1513:

A=L*W Let w be width L=(w+6) 40=(w+6)*w 40= =0 w(w+10)-4(w+10)=0 (w-4)(w+10)=0 w=4 Length= 10 units Width=4 units (so it make it 10 by 4)

DogGirl27:

thank youu

Alex1513:

np